# Statistical significance in choice modelling: computation, usage and reporting

### v2

@@ 6 Discussion and conclusions
This note has looked at the issues of the interpretation and reporting of measures of statistical confidence in the context of choice model estimation. The misuse of the notion of “statistical significance” has been a long-standing concern for statisticians in particular and science more widely. The authoritative paper by Wasserstein et al. (2019) goes as far as concluding “that it is time to stop using the term ‘statistically significant’ entirely. Nor should variants such as ‘significantly different’, ‘p<0.05’, and ‘nonsignificant’ survive, whether expressed in words, by asterisks in a table, or in some other way.” Less controversially, they say that “[analysts should not] believe that an association or effect exists just because it was statistically significant [or] that an association or effect is absent just because it was not statistically significant.”
While we do not foresee a situation where choice modellers will abandon the notion completely, at least any time soon, we are of the opinion that a more nuanced approach is required. We believe that:
@@
  • Analysts should look further than statistical significance alone. In a health context, it is common to talk about “clinical importance”, which measures whether a treatment has a genuine and noticeable effect on health outcomes. In choice modelling more broadly, analysts may wish to consider the “behavioural importance” of a parameter, i.e. whether it changes predictions (cf. discussions in Section 4.1.2 and the “policy importance”, i.e. whether a finding has a sizeable impact on the outcome of any process using the results).
@@
  • Analysts need to properly document their test findings, by clarifying whether these are one-sided or two-sided.
@@
Finally, and as an alternative to frequentist significance testing, analysts could consider the Bayesian approach to parameter uncertainty, where prior knowledge is updated in response to evidence provided by data, and statistical inference is based upon decision-making theory. Through the concept of posterior probabilities, Bayesian econometrics offers a probabilistic representation of not only parameters but also hypotheses. For instance, Bayes estimates are the whole posterior distribution of the parameters from which analysts can derive the exact probability of (credible) intervals containing the true parameter. Bayes’ rule can also be used to derive measures of relative evidence across competing models, including posterior probabilities of hypotheses. For example, and unlike $p$-values, Bayes factors provide a data-supported measure of the odds in favour of the null hypothesis over the alternative. Even though Bayesian hypothesis testing is more intuitive and departs from frequentist concepts, there is an open discussion among Bayesian practitioners about the flaws in null hypothesis significance testing, which goes beyond the scope of this note.
@@ Acknowledgements
Stephane Hess acknowledges the support of the European Research Council through the advanced grant 101020940-SYNERGY. Angelo Guevara acknowledges the support by ANID FONDECYT 1231584 and ANID PIA/PUENTE AFB220003.

### v3

@@ 6 Discussion and conclusions
This note has looked at the issues of the interpretation and reporting of measures of statistical confidence in the context of choice model estimation. We have provide an in-depth overview of the reason for parameter uncertainty, its computation, and its use in confidence intervals and hypothesis testing. The misuse of the notion of “statistical significance” has been a long-standing concern for statisticians in particular and science more widely. The authoritative paper by Wasserstein et al. (2019) goes as far as concluding “that it is time to stop using the term ‘statistically significant’ entirely. Nor should variants such as ‘significantly different’, ‘p<0.05’, and ‘nonsignificant’ survive, whether expressed in words, by asterisks in a table, or in some other way.” Less controversially, they say that “[analysts should not] believe that an association or effect exists just because it was statistically significant [or] that an association or effect is absent just because it was not statistically significant.”
We share the view that excessive attention is paid to the statistical properties of findings, often at the expense of the behavioural and policy implications. At the same time, we do not foresee a situation where choice modellers will abandon the notion completely, at least any time soon, and this note thus provides some guidance for applied work. We believe that:
@@
  • Analysts should look further than statistical significance alone. In a health context, it is common to talk about “clinical importance”, which measures whether a treatment has a genuine and noticeable effect on health outcomes. In choice modelling more broadly, analysts may wish to consider the “behavioural importance” of a parameter, i.e. whether it changes predictions (cf. discussions in Section 4.3 and the “policy importance”, i.e. whether a finding has a sizeable impact on the outcome of any process using the results).
@@
  • Analysts need to properly document their test findings, by clarifying whether these are for one-sided or two-sided tests.
@@
An issue that deserves some discussion at the end is that of researcher incentives and the related problem of $p$-hacking, which essentially implies coming up with a specification that gives an analyst the right $t$-ratio for the effect they’re interested in. Choice modelling is largely an empirical field, and it is noticeable how in some topic areas, a lot of emphasis is placed on a priori registration of hypotheses, while in other areas, a more data-driven approach is used. While the latter is often described as fishing for evidence (or significance), it should also be clear that the former carries risks too, namely in terms of findings being biased by effects that are unaccounted for. Furthermore, analysts wary of the straitjacket of pre-registration may be tempted to develop a near-complete list of possible hypotheses, which defeats the entire point. However, such bad practice is unlikely to change without better editorial practice and improved teaching of statistics, not just for authors but also for reviewers who continuously call for tables of results filled with stars and only significant parameters.
An additional caveat relates to the disconnect between applied choice modelling practice and inference theory. The latter is based on the idea that the model is fixed. In practice, we conduct tests on individual parameters for one model at a time, and these tests only tell us something subject to the hypothesis that this is the true model. A different conclusion could be reached with a different model. Different $p$-values will be obtained for the same parameter in different models, which relates to the earlier point that the wrongly specified model has not only wrong estimates but also wrong statistical significance properties.
Finally, and as an alternative to frequentist significance testing, analysts could consider the Bayesian approach to parameter uncertainty, where prior knowledge is updated in response to evidence provided by data, and statistical inference is based upon decision-making theory. Through the concept of posterior probabilities, Bayesian econometrics offers a probabilistic representation of not only parameters but also hypotheses. For instance, Bayes estimates are the whole posterior distribution of the parameters from which analysts can derive the exact probability of (credible) intervals containing the true parameter. Bayes’ rule can also be used to derive measures of relative evidence across competing models, including posterior probabilities of hypotheses. For example, and unlike $p$-values, Bayes factors provide a data-supported measure of the odds in favour of the null hypothesis over the alternative. Even though Bayesian hypothesis testing is more intuitive and departs from frequentist concepts, there is an open discussion among Bayesian practitioners about the flaws in null hypothesis significance testing, which goes beyond the scope of this note, and we refer the reader to McElreath (2020).
@@ Acknowledgements
Stephane Hess acknowledges the support of the European Research Council through the advanced grant 101020940-SYNERGY. Angelo Guevara acknowledges the support by ANID FONDECYT 1231584 and ANID PIA/PUENTE AFB220003. The authors would like to thank David Bunch for feedback on earlier work. We are also grateful to two anonymous referees whose comments substantially improved the paper.